OpenStudy (anonymous):

If 2tanx/1-tan^2 x = then x can equal:? there can be more than 1 answer x= 7pi/8 + npi x= 5pi/8 + npi x= 3pi/8 + npi x= pi/8 + npi

OpenStudy (aggie12341):

yes

OpenStudy (aggie12341):

x=pi/8+npi

OpenStudy (solomonzelman):

I don;t get the question (iether)

OpenStudy (aggie12341):

add me solomonzelman be my fan

OpenStudy (primeralph):

@AlvaroBravo Question not complete.

OpenStudy (solomonzelman):

is x/1-tan^2 x \[\frac{2tanx}{1-\tan^2x}\] or \[\frac{2tanx}{1} - \tan^2x\]

OpenStudy (anonymous):

the first one